OpenStudy (4itc):

if x:y = 5:2 and y:z = 3:2, what is the ratio of x:z ? a) 3:1 b) 3:5 c) 5:3 d) 8:4 E) 15:4

OpenStudy (4itc):

i got 5:2...... yea.

Vocaloid (vocaloid):

good guess but not quite

Vocaloid (vocaloid):

an easy way to solve this is to pick numbers for x, y, and z

Vocaloid (vocaloid):

if x : y = 5 : 2 we can let x = 10 and y = 4

Vocaloid (vocaloid):

so, if y = 4 and the ratio between y and z is y: z = 3 : 2 what is z?

OpenStudy (4itc):

so you times it by two!? each number that is. z is 4.

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

y: z = 3 : 2 means y = (3/2)z

Vocaloid (vocaloid):

so if y = 4, what is z?

OpenStudy (4itc):

3/2 = 1.5 = z?

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

4 = (3/2)z what is z?

OpenStudy (4itc):

3/2 x 4/1 = 12/2 = 6...... but thad be 6z... 2/3 x 4/1 = 8/3 = 2.6666666667. grrr ill go with 6

Vocaloid (vocaloid):

good, we can just leave it at 8/3

Vocaloid (vocaloid):

now we can write our new ratio

Vocaloid (vocaloid):

x : z = 10 : 8/3 we can multiply both x and z by 3 to get 30 : 8 then we can divide both by 2 to get 15 : 4

Vocaloid (vocaloid):

so E is your answer

Vocaloid (vocaloid):

also, if you're wondering how I got x = 10 and y = 4 we could have used any x and y values that fit the ratio 5 : 2

Vocaloid (vocaloid):

so, we could have used x = 5 and y = 2 or x = 15 and y = 6, etc...

OpenStudy (4itc):

thanks! you helped a lot!